# Holographic analysis of boundary correlation functions for the hyperbolic-lattice Ising model

### v1

@@ 1 Introduction
Figure 1: A $\{5,4\}$ hyperbolic-lattice Ising model illustrated on the Poincarè unit disk, where all regular pentagons are equivalent in the hyperbolic space. The circles on the lattice nodes represent Ising spins; Nodes near the outer edge of the unit disk are omitted for simplicity. The generation index $n$ for the nodes along the horizontal geodesic line corresponds to the hyperbolic distance from the center node, and $N$ denotes the generation index for the outer edge node. One of the four quadrants sharing the center node (shaded region) corresponds to a CTM $A$.
@@ 2.1 $\{5,4\}$ lattice model and CTMRG
where $\sigma_{i}=\pm 1$ denotes an Ising spin variable at the $i$th node on the $\{5,4\}$ lattice, and $\langle i,j\rangle$ represents the sum with respect to the nearest neighboring pairs. $K\equiv\beta J$ denotes an inverse temperature with the exchange coupling $J=1$. Also, $h_{b}$ represents a magnetic field applied to spins at the outer edge boundary, and thus, no magnetic field exists in the bulk region. Moreover, we have introduced an effective coordination number for edge spins as $\mu=1$ or $2$, taking into account the number of bonds connected with edge spins. In Fig. 3, for example, $\mu=1$ for $k=0$, 2, 5, 7, 10 and 12, while $\mu=2$ for $k=1$, 3, 4, 6, 8, 9, and 11. We assume the free boundary condition for the outer edge boundary spins.
@@
As in Figs. 1, the hyperbolic lattice is usually drawn on Poincarè unit disk. Here, we briefly summarize the geometric property of the hyperbolic tiling with a regular polygon whose edge length is normalized to be unity. As in Appendix A, the hyperbolic distance for the Poincarè unit disk coordinate is generally defined by the metric (21) with the Poincarè radius $L$. For the polygon tiling on the hyperbolic plane, a crucial point is that the shape of a polygon defines a particular value of $L$; Let us write the radius for the $\{p,q\}$ tiling as $L_{pq}$. Then, the hyperbolic version of the cosine formula provides
@@
for the $\{5,4\}$ tiling. At Poincarè unit disk level, an approximated radius $r_{n}$ associated with the $n$th generation circle is given by
@@
In turn, we can also introduce another effective coordinate for the hyperbolic tiling lattice based on the number of nodes along its outer edge circle, different from the above standard definition of $L_{pq}$. Let us consider lattice nodes with a generation index $n$, which are approximately distributed on a circle defined by a hyperbolic distance $n$ from the center of the disk. As in Appendix A, in general, the circumference of the circle with a hyperbolic distance $\rho$ in the hyperbolic plane with a Poincarè radius $L$ is given by $2\pi L\sinh\left(\frac{\rho}{L}\right)\sim\exp\left(\frac{\rho}{L}\right)$ for $\rho\gg L$ [See Eq. (27)]. Assuming a uniform distribution of the nodes with the $n$th generation in the circumference direction, we may approximately write the $n$ dependence of the number of lattice nodes as
@@
with $p_{\mathrm{eff}}=3.73\cdots$ from Table 1.Serina2016 By comparing Eqs. (6) and (7), we can estimate the effective Poincarè radius as
@@
which slightly deviates from Eq. (4). As in the case of the Bethe lattice, then, we may describe the distribution of nodes by introducing an effective coordinate corresponding to Poincarè upper half (37) with compactification, summarized in Appendix A.2.  [^2] For the Bethe lattice case in Ref. Oku2023 , all the edge nodes of the tree network were naively arranged in the outer circle of the disk-like representation, for which the effective metric near the outer edge boundary corresponds to (compactified) Poincarè upper-half plane, although the appearance of the tree network looks as if Poincarè unit disk. Here, it may be interesting to note that $L_{\mathrm{eff}}$ is based on the macroscopic property (6) for the hyperbolic plane. In contrast, $L_{pq}$ purely originates from the microscopic structure of the regular polygon tiling.
@@ 3 Correlation functions for the $\{5,4\}$ lattice model
Let us precisely analyze the correlation function in Fig. 5, exhibiting power-law decay with quasi-periodic oscillation. First, we note that the quasi-periodic oscillation is a general feature of the hyperbolic tiling because the arrangement of the $P$ and $A$ along the circumference direction has quasi-periodicity. Such quasi-periodicity for the boundary system is also discussed in terms of the matchgate tensor networkJahn2022 and the holographic inflation ruleBasteiro2022 . Then, the resulting stable power-law behaviors of the correlation function suggests that the quasi-periodic oscillation originates from the microscopic details around each node and is irrelevant in analyzing their scaling dimensions.
@@ 5 Cutoff effect of the bond dimension and renormalization group
As demonstrated in Fig. 8(b), the correlation functions across the center spin of the CTM are actually described well even if the cutoff is introduced, but those for $\tilde{\rho}=1\sim 8$ rapidly loses accuracy. In the entanglement context, another important observation for the hyperbolic-lattice model is that the decay of the CTM spectrum is very rapid even at the bulk critical point $K_{c}$, which is consistent with the very short correlation length at $K_{c}$ in the bulk region (See $\xi$ in Fig. 4). Thus, a few dominant eigenvalues of the CTM is enough to represent the particular long-range entanglement across the center node. Then, the hyperbolic-lattice geometry transmits the high accuracy along the CTM edges into those of the boundary spins near the two edges. As a result, the CTMRG nature incorporated with the hyperbolic-lattice geometry realizes a selective RG transformation of long-distance correlations. On the other hand, a massive superposition of the tiny eigenvalue eigenvectors of the CTM is necessary for reproducing the short-range correlation for the outer edge within the left/right block.
From the holographic entanglement point of view, verifying the Ryu-Takayanagi formula for a boundary state defined on the outer edge of hyperbolic-lattice networks is highly desired. However, we should note that, unlike the MERA-type algorithmsMERA2007 ; TNR2015 , the theoretical framework of CTMRG is not appropriate for directly computing such an entanglement entropy for the outer edge boundary.
@@ 6 Summary and discussions
We have precisely investigated the boundary spin correlation function for the hyperbolic-lattice Ising model and its relation to the bulk correlation length along the geodesic line. The CTMRG computation for the $\{5,4\}$ lattice model has demonstrated that in the high-temperature (disorder) region, the boundary correlation function exhibits power-law decay with quasi-periodic oscillation, while the bulk correlation function always exhibits exponential decay. Taking account of the geometric relation of the distance along the outer edge boundary and the bulk correlation path based on the geodesic line, we have found that the most relevant part of the regular peak-like structure in the oscillating correlation function can be characterized by the scaling dimension $\Delta_{\mathrm{eff}}$ originating from the radius $L_{\mathrm{eff}}$ for the effective Poincarè upper-half based on the exponential inflation of boundary nodes. On the other hand, we have also revealed that the dominant center line in the quasi-periodically oscillating behavior is explained well by the scaling dimension $\Delta_{54}$ associated with the radius $L_{54}$ for the Poincarè unit disk for the $\{5,4\}$ tiling network. In the low-temperature (bulk ferromagnetic order) region, moreover, we have demonstrated the crossover behavior of the boundary correlation function, depending on the depth of the corresponding correlation path. Then, the scaling dimensions for the two crossover regime can be also well described by the bulk correlation lengths for the unstable disorder and bulk ferromagnetic fixed points.
@@
From the tensor network point of view, we have pointed out that the CTMRG combined with the hyperbolic lattice maintains the particular long-range correlation for the boundary spins near the CTM edges. Although this behavior is similar to the Kadanoff-Wilson type real-space RGWilsonRG ; Kadanoff2014 , the mechanism of such a selective RG transformation in CTMRG can be explained in analogy with the MPS for the linear chain. However, the current CTMRG algorithm is insufficient to reveal entanglement entropies associated with outer edge boundary spins in the hyperbolic lattice. Therefore, it is also intriguing to elucidate to what extent the MPS is able to capture the entanglement structure under the hyperbolic geometry.
@@ Appendix A Unit disk representations for the hyperbolic plane
A.1 Poincarè unit disk
The metric for Poincarè unit disk for the hyperbolic plane is given by
@@ A.1 Poincarè unit disk
By regarding $\rho$ as a network distance from the center node $n$, the corresponding radius in Poincarè unit disk is
@@
Figure 9: A geodesic line for the Poincarè unit disk. $t$ denotes an angle parameter.
@@
A.2 An effective Poincarè upper-half plane
@@
where $L_{\mathrm{eff}}=1/\log p_{\mathrm{eff}}$ and $x$ is rescaled by $\tilde{x}\equiv\frac{1}{2\pi\kappa\log p_{\mathrm{eff}}}$, as in the case of the Bethe lattice model. This coordinate corresponds to the effective Poincarè upper half with compactification for $\tilde{x}$ due to the periodic boundary condition. The geodesic length for a path connecting two points, P and Q, on the outer edge boundary is given by
@@
Here, it should be emphasized that Eq. (37) is not directly related to Eq. (21) by the standard conformal mapping ($w=\frac{z-i}{z+i}$). As the geodesic line comes into the deep interior of the unit disk, thus, Eq. (38) deviates from the geodesic length of Eq. (33) for the Poincarè unit disk. Here, note that the definition of the cutoff scale in Eq. (38) does not correspond to $a$ in Eq.(38). Also, Eq. (37) is valid only for $z\ll 1$ while Eq. (21) is able to describe Porencare’s unit disk correctly.

### v2

@@ 1 Introduction
Figure 1: A $\{5,4\}$ hyperbolic-lattice Ising model illustrated on the Poincaré unit disk, where all regular pentagons are equivalent in the hyperbolic space. The circles on the lattice nodes represent Ising spins; Nodes near the outer edge of the unit disk are omitted for simplicity. The generation index $n$ for the nodes along the horizontal geodesic line corresponds to the hyperbolic distance from the center node, and $N$ denotes the generation index for the outer edge node. One of the four quadrants sharing the center node (shaded region) corresponds to a CTM $A$.
@@ 2.1 $\{5,4\}$ lattice model and CTMRG
where $\sigma_{i}=\pm 1$ denotes an Ising spin variable at the $i$th node on the $\{5,4\}$ lattice, and $\langle i,j\rangle$ represents the sum with respect to the nearest neighboring pairs. $K\equiv\beta J$ denotes an inverse temperature with the exchange coupling $J=1$. Also, $h_{b}$ represents a magnetic field applied to spins at the outer edge boundary, and thus, no magnetic field exists in the bulk region. Moreover, we have introduced an effective coordination number for edge spins as $\mu=1$ or $2$, taking into account the number of bonds connected with edge spins; This is because, if we think of a holographic renormalization of spins from the outer side, the resulting boundary magnetic fields depend chiefly on the number of bonds linked to the outer spins traced out. In Fig. 3, for example, $\mu=1$ for $k=0$, 2, 5, 7, 10, and 12, while $\mu=2$ for $k=1$, 3, 4, 6, 8, 9, and 11. We assume the free boundary condition for the outer edge boundary spins.
@@
As in Figs. 1, the hyperbolic lattice is usually drawn on Poincaré unit disk. Here, we briefly summarize the geometric property of the hyperbolic tiling with a regular polygon whose edge length is normalized to be unity. As in Appendix A, the hyperbolic distance for the Poincaré unit disk coordinate is generally defined by the metric (21) with the Poincaré radius $L$. For the polygon tiling on the hyperbolic plane, a crucial point is that the shape of a polygon defines a particular value of $L$; Let us write the radius for the $\{p,q\}$ tiling as $L_{pq}$. Then, the hyperbolic version of the cosine formula provides
@@
for the $\{5,4\}$ tiling. At Poincaré unit disk level, an approximated radius $r_{n}$ associated with the $n$th generation circle is given by
@@
In turn, we can also introduce another effective coordinate for the hyperbolic tiling lattice based on the number of nodes along its outer edge circle, different from the above standard definition of $L_{pq}$. Let us consider lattice nodes with a generation index $n$, which are approximately distributed on a circle defined by a hyperbolic distance $n$ from the center of the disk. As in Appendix A, in general, the circumference of the circle with a hyperbolic distance $\rho$ in the hyperbolic plane with a Poincaré radius $L$ is given by $2\pi L\sinh\left(\frac{\rho}{L}\right)\sim\exp\left(\frac{\rho}{L}\right)$ for $\rho\gg L$ [See Eq. (27)]. Assuming a uniform distribution of the nodes with the $n$th generation in the circumference direction, we may approximately write the $n$ dependence of the number of lattice nodes as
@@
with $p_{\mathrm{eff}}=3.73\cdots$ from Table 1.Serina2016 By comparing Eqs. (6) and (7), we can estimate the effective Poincaré radius as
@@
which clearly deviates from Eq. (4). As in the case of the Bethe lattice, then, we may describe the distribution of nodes by introducing an effective coordinate corresponding to Poincaré upper half (37) with compactification, summarized in Appendix A.2.  [^2] For the Bethe lattice case in Ref. Oku2023 , all the edge nodes of the tree network were naively arranged in the outer circle of the disk-like representation, for which the effective metric near the outer edge boundary corresponds to (compactified) Poincaré upper-half plane, although the appearance of the tree network looks as if Poincaré unit disk. Here, it may be interesting to note that $L_{\mathrm{eff}}$ is based on the macroscopic property (6) for the hyperbolic plane. In contrast, $L_{pq}$ purely originates from the microscopic structure of the regular polygon tiling.
@@ 3 Correlation functions for the $\{5,4\}$ lattice model
Let us precisely analyze the correlation function in Fig. 5, exhibiting power-law decay with quasi-periodic oscillation. First, we note that the quasi-periodic oscillation is a general feature of the hyperbolic tiling because the arrangement of the $P$ and $A$ along the circumference direction has quasi-periodicity. Such quasi-periodicity for the boundary system is also discussed in terms of the matchgate tensor networkJahn2022 and the holographic inflation ruleBasteiro2022 ; Basteiro2023 . Then, the resulting stable power-law behaviors of the correlation function suggests that the quasi-periodic oscillation originates from the microscopic details around each node and is irrelevant in analyzing their scaling dimensions.
@@ 5 Cutoff effect of the bond dimension and renormalization group
As demonstrated in Fig. 8(b), the correlation functions across the center spin of the CTM are actually described well even if the cutoff is introduced, but those for $\tilde{\rho}=1\sim 8$ rapidly loses accuracy. In the context of the CTM-based entanglement, another important observation for the hyperbolic-lattice model is that the decay of the CTM spectrum is very rapid even at the bulk critical point $K_{c}$, which is consistent with the very short correlation length at $K_{c}$ in the bulk region (See $\xi$ in Fig. 4). Thus, a few dominant eigenvalues of the CTM is enough to represent the particular long-range entanglement across the center node. Then, the hyperbolic-lattice geometry transmits the high accuracy along the CTM edges into those of the boundary spins near the two edges. As a result, the CTMRG nature incorporated with the hyperbolic-lattice geometry realizes a selective RG transformation of long-distance correlations. On the other hand, a massive superposition of the tiny eigenvalue eigenvectors of the CTM is necessary for reproducing the short-range correlation for the outer edge within the left/right block.
From the holographic entanglement point of view, verifying the Ryu-Takayanagi formula for a boundary state defined on the outer edge of hyperbolic-lattice networks is highly desired. However, we should note that if the hyperbolic lattice Ising model corresponds to a counterpart system on the AdS side, as in the present case, an appropriate definition of entanglement entropy at the level of the tensor network framework remains a nontrivial problem. On the other hand, if the boundary state of the hyperbolic lattice system is regarded as a wavefunction for a boundary CFT, it may be possible to use the MERA-type algorithmsMERA2007 ; TNR2015 for computing the corresponding entanglement entropy.
@@ 6 Summary and discussions
We have precisely investigated the boundary spin correlation function for the hyperbolic-lattice Ising model and its relation to the bulk correlation length along the geodesic line. The CTMRG computation for the $\{5,4\}$ lattice model has demonstrated that in the high-temperature (disorder) region, the boundary correlation function exhibits power-law decay with quasi-periodic oscillation, while the bulk correlation function always exhibits exponential decay. Taking account of the geometric relation of the distance along the outer edge boundary and the bulk correlation path based on the geodesic line, we have found that the most relevant part of the regular peak-like structure in the oscillating correlation function can be characterized by the scaling dimension $\Delta_{\mathrm{eff}}$ originating from the radius $L_{\mathrm{eff}}$ for the effective Poincaré upper-half based on the exponential inflation of boundary nodes. On the other hand, we have also revealed that the dominant center line in the quasi-periodically oscillating behavior is explained well by the scaling dimension $\Delta_{54}$ associated with the radius $L_{54}$ for the Poincaré unit disk for the $\{5,4\}$ tiling network. In the low-temperature (bulk ferromagnetic order) region, moreover, we have demonstrated the crossover behavior of the boundary correlation function, depending on the depth of the corresponding correlation path. Then, the scaling dimensions for the two crossover regime can be also well described by the bulk correlation lengths for the unstable disorder and bulk ferromagnetic fixed points.
@@
From the tensor network point of view, we have pointed out that the CTMRG combined with the hyperbolic lattice maintains the particular long-range correlation for the boundary spins near the CTM edges. Although this behavior is similar to the Kadanoff-Wilson type real-space RGWilsonRG ; Kadanoff2014 , the mechanism of such a selective RG transformation in CTMRG can be explained in analogy with the MPS for the linear chain. In this sense, it is intriguing to elucidate to what extent the MPS is able to resolve the entanglement structures under the hyperbolic geometry. However, we should also note that tensor networks such as CTMRG are insufficient for the holographic version of the entanglement entropy associated with outer edge boundary spins in the hyperbolic lattice Ising model.
@@ Appendix A Unit disk representations for the hyperbolic plane
A.1 Poincaré unit disk
The metric for Poincaré unit disk for the hyperbolic plane is given by
@@ A.1 Poincaré unit disk
By regarding $\rho$ as a network distance from the center node $n$, the corresponding radius in Poincaré unit disk is
@@
Figure 9: A geodesic line for the Poincaré unit disk. $t$ denotes an angle parameter.
@@
A.2 An effective Poincaré upper-half plane
@@
where $L_{\mathrm{eff}}=1/\log p_{\mathrm{eff}}$ and $x$ is rescaled by $\tilde{x}\equiv\frac{1}{2\pi\kappa\log p_{\mathrm{eff}}}$, as in the case of the Bethe lattice model. This coordinate corresponds to the effective Poincaré upper half with compactification for $\tilde{x}$ due to the periodic boundary condition. The geodesic length for a path connecting two points, P and Q, on the outer edge boundary is given by
@@
Here, it should be emphasized that Eq. (37) is not directly related to Eq. (21) by the standard conformal mapping ($w=\frac{z-i}{z+i}$). As the geodesic line comes into the deep interior of the unit disk, thus, Eq. (38) deviates from the geodesic length of Eq. (33) for the Poincaré unit disk. Here, note that the definition of the cutoff scale in Eq. (38) does not correspond to $a$ in Eq.(38). Also, Eq. (37) is valid only for $z\ll 1$ while Eq. (21) is able to describe Porencare’s unit disk correctly.